b) You visit your Manufacturing division, which has a machine producing metal bolts. In a tray of these bolts, 94% are within the allowable diameter tolerance value. The remainder exceed the tolerance. You withdraw six bolts at random from the tray. Determine the probabilities that;

(i) Two of the six bolts exceed the diameter.

(ii) More than two of the six bolts exceed the diameter.
Answer #1

Here P(Witning allowable diameter tolerance value) = p = 0.94

P(Not within allowable diameter tolerance value) = q = 0.06

Here number of bolts = n = 6

(i) P(x = 2) = BINOMIAL (x = 2; n = 6; q = 0.06) = 6C2 (0.06)2(0.94)4  = 0.0422

(ii) P(x > 2) = 1 - P(x 2) = 1 - BINOMIAL (2; 6 ; 0.06; true) = 1 - 0.9962 = 0.0038